OpenStudy (anonymous):

x + x + 2 + 2 + 2 2(x + 3) 2x + 23 4 + x + 2 + x Question: a. Which of the given expressions is NOT equivalent to the expression 2x + 6? Why? If you were to correct it, what should it be?

OpenStudy (anonymous):

Write an expression equivalent to 4x + 12 using the distributive property. c. Write an expression equivalent to 4x + 12 by combining like terms.

OpenStudy (anonymous):

these r 3 questions :)

OpenStudy (anonymous):

i know this :D

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

a.) If you simplify which of the problems does not equal 2x + 6

OpenStudy (anonymous):

could you i was thinking the second answer

OpenStudy (anonymous):

x + x + 2 + 2 + 2 x+x = 2x 2+2+2=6 2x + 6

OpenStudy (wmj259):

@heyyitspaige, In my honest opinion paige, i deep down know that you know how to do this.

OpenStudy (anonymous):

uhhh forget about the could you i wrote was a mistake :D

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

2(x + 3) (Use distributive property) (2x)+(2*3) 2x + 6

OpenStudy (anonymous):

omg its actually the 3rd one rirght?

OpenStudy (anonymous):

2x + 23 (I'm guessing you mean 2x + 2*3?)

OpenStudy (anonymous):

so what should it be it says if u were to correct it.. what should it be?

OpenStudy (anonymous):

You mean the way i wrote it or the way your wrote it? (the way your wrote it it's as simplified as can be, the way i wrote it it simplifies to 2x + 6)

OpenStudy (anonymous):

''Which of the given expressions is NOT equivalent to the expression 2x + 6? Why? If you were to correct it, what should it be?''

ganeshie8 (ganeshie8):

I think 2x + 23 does not equal 2x+6 so that should be the answer

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Well wait Check the last one @ganeshie8

OpenStudy (anonymous):

the last what? XD

OpenStudy (anonymous):

4 + x + 2 + x = 6 + x

ganeshie8 (ganeshie8):

last one : 4 + x + 2 + x 4 + 2 + x + x 6 + 2x 2x + 6 this is fine right ?

OpenStudy (anonymous):

6 + 2x*

OpenStudy (anonymous):

Or 2x + 6

OpenStudy (anonymous):

So it is indeed 3

OpenStudy (anonymous):

ok im confused :/

OpenStudy (anonymous):

i know its the 3rd one...

OpenStudy (anonymous):

Why? If you were to correct it, what should it be? Because 2x + 23 does not have any relation to 2x + 6. To correct it add a multiplication symbol and solve: 2x + 2*3 = 2x + 6

OpenStudy (anonymous):

what is 2x+23?

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

youve been so much help!

OpenStudy (anonymous):

Write an expression equivalent to 4x + 12 using the distributive property. So what are both 4x and 12 divisible by?

OpenStudy (anonymous):

@heyyitspaige ?

OpenStudy (anonymous):

oh sry im writing the answer for the first one ;)

OpenStudy (anonymous):

lol no problem

OpenStudy (anonymous):

kk next question :)

OpenStudy (anonymous):

Write an expression equivalent to 4x + 12 using the distributive property. So what are both 4x and 12 divisible by?

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

im so bad at math so sry :( :(

OpenStudy (anonymous):

Divisible by not factors of lol

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

ughhh i feel so dumb

OpenStudy (anonymous):

You're ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

4 adn 12 are both divisible by 4

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So divide 4 from both 4x and 12

OpenStudy (anonymous):

Now you have x + 3

OpenStudy (anonymous):

kk :)

OpenStudy (anonymous):

Add 4 as a factor to the beginning so now you have the answer of 4(x+3)

OpenStudy (anonymous):

thank you! again :/

OpenStudy (anonymous):

lol it's ok

OpenStudy (anonymous):

im so lazy :o yawns*

OpenStudy (anonymous):

Now to complete that you may have to put (4x)+(4*3)..... but I'm not positive

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

so were on the 3rd question now?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Write an expression equivalent to 4x + 12 by combining like terms.

OpenStudy (anonymous):

But all like terms are already combined.... @ganeshie8 can you help out please?

OpenStudy (anonymous):

yea... thats confusing me sorta

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

@Ashleyisakitty Can you help us out here please?

OpenStudy (anonymous):

@Ashleyisakitty Can you help us out here please?

OpenStudy (anonymous):

Sorry i have to go now... good luck!

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

i need help... last question! :/

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@nincompoop